OpenStudy (anonymous):

jeanine paid $599.20 for a TV that was discounted 20%. what was the original price of the TV? please help show ALL work problem.

pooja195 (pooja195):

If it was 20% discounted, then it costed 80% of the original price, so 599.20 = 0.8 * original price original price = 599.2/0.8= 749.

OpenStudy (anonymous):

Let original price be $x, since she got 20% discount, you have 80% of x = 599.20 x = 599.20/0.8

OpenStudy (zehanz):

Or, because 599.20 is 80%, first divide 599.20 by 80 to get 1%. Then multiply that with 100 to get the original price.

OpenStudy (anonymous):

Thanks guys. :)

OpenStudy (zehanz):

YW!

pooja195 (pooja195):

@purple99 by the way WELCOME TO THIS WEBSITE!!!

pooja195 (pooja195):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Cool.

pooja195 (pooja195):

ok first people with the purple A are ambassador u can ask them anything you want in messages of course las don't SPAM ENJOY!!!!

OpenStudy (anonymous):

Ok